# Avoiding Premature Collapse: Adaptive Annealing for Entropy-Regularized Structural Inference

### v1

@@ Abstract
Differentiable matching layers, often implemented via entropy-regularized Optimal Transport, serve as a critical approximate inference mechanism in structural prediction. However, recovering discrete permutations via annealing $\epsilon\to 0$ is notoriously unstable. We identify a fundamental mechanism for this failure: Premature Mode Collapse. By analyzing the non-normal dynamics of the Sinkhorn fixed-point map, we reveal a theoretical thermodynamic speed limit. Under standard exponential cooling, the shift in the target posterior ($O(1)$) outpaces the contraction rate of the inference operator, which degrades as $O(1/\epsilon)$. This mismatch inevitably forces the inference trajectory into spurious local basins. To address this, we propose Efficient PH-ASC, an adaptive scheduling algorithm that monitors the stability of the inference process. By enforcing a linear stability law, we decouple expensive spectral diagnostics from the training loop, reducing overhead from $O(N^{3})$ to amortized $O(1)$. Our implementation and interactive demo are available at https://github.com/xxx0438/torch-sinkhorn-asc and https://huggingface.co/spaces/leon0923/torch-sinkhorn-asc-demo. bounded away from zero in generic training dynamics unless the feature extractor converges unrealistically fast.
@@ 3.1 Problem Setup and Sinkhorn Dynamics
See Appendix LABEL:app:proofs_sensitivity for sensitivity analysis and Appendix LABEL:app:proofs_spectral for spectral bounds. ∎
@@ 3.2 The Adiabatic Tracking Theorem
The proofs for Theorem 3.2 and Corollary 3.3 are provided in Appendix B.
@@ A.5 Directional sensitivity lower bound
A.6 Implicit Sensitivity Lower Bound
@@
Under the Non-Degeneracy Assumption A.1, the sensitivity of the Sinkhorn fixed point with respect to the cost matrix scales inversely with the temperature:
$DS_{\epsilon}(C)||_{op}\geq\frac{c_{0}}{\epsilon}$ (17)
for some constant $c_{0}>0$ that depends on the connectivity of the active support but is independent of $\epsilon$ for sufficiently small $\epsilon$.
@@
We utilize the implicit function theorem on the active support set $S$. The Sinkhorn map on the active set can be structurally written as a smooth function $F(P,C,\epsilon)=0$. Specifically, the mapping from cost $C$ to transport plan $P$ involves the exponential term $\exp(-C_{ij}/\epsilon)$. Let us consider a perturbation along a specific direction in the cost space. Specifically, consider the directional derivative along the active components. The relationship between the primal perturbation $\Delta P$ and cost perturbation $\Delta C$ on the active subspace is given by the linear system derived in Appendix A.3:
$\mathcal{A}_{S}(\epsilon)\begin{pmatrix}\dot{f}\\ \dot{g}\end{pmatrix}=\frac{1}{\epsilon}\mathcal{B}_{S}(\Delta C)$ (18)
Crucially, the operator $\mathcal{A}_{S}(\epsilon)$ converges to a finite, non-singular limit matrix $\mathcal{A}_{0}$ determined by the bipartite graph Laplacian of the active support as $\epsilon\to 0$ (modulo the gauge constraints). Let $M(\epsilon)=R_{S}\mathcal{A}_{S}(\epsilon)^{-1}\mathcal{B}_{S}$ be the effective linear map such that $\Delta P=\frac{1}{\epsilon}M(\epsilon)\Delta C$. The matrix $M(\epsilon)$ captures the geometric response of the polytope vertices to cost changes. Since the active support $S$ corresponds to a stable matching (Assumption A.1), the structural response matrix $M(\epsilon)$ is non-vanishing. Specifically, there exists at least one direction $\Delta C^{*}$ (e.g., increasing the cost of a single active edge) that forces a mass redistribution. For this direction:
[…]
Since $M(\epsilon)$ depends only on the limit marginals and support structure, $\lim_{\epsilon\to 0}\sigma_{\min}(M(\epsilon))=\mu>0$. Thus, the operator norm satisfies:
[…]
This confirms the $O(1/\epsilon)$ scaling behavior. ∎
Remark on Numerical Verification.
While the theoretical lower bound relies on the structural non-degeneracy $\mu>0$, in Section 6 we empirically verify this by performing Singular Value Decomposition (SVD) on the Jacobian $J_{\epsilon}$ during training. We consistently observe that the largest singular values of the sensitivity operator scale strictly as $\epsilon^{-1}$, validating the lemma.
@@ A.9 Raw numeric estimates and CSV
[…]
@@ B.2 Lemma: Equilibrium Error Bound
Non-normal transient amplification despite spectral contraction Theorem LABEL:thm:non_normal_stability; full proof in Appendix LABEL:app:proof_non_normal
Effective basin shrinkage by modal condition number $\kappa(V)$ Appendix LABEL:app:proof_non_normal, transient growth bound
Discrete-time tracking formulation for annealing Appendix LABEL:app:tracking_full, Section “Discrete Tracking Model”
Existence of a dynamical annealing speed limit Theorem 3.2; proof and constants in Appendix LABEL:app:tracking_full
@@
Justification of QSA drift estimation via implicit differentiation Appendix A.3; error control in Appendix LABEL:app:tracking_full
ASC admissible drift threshold $\|\Delta_{t}\|\leq\frac{1-\rho(J_{\varepsilon})}{\kappa(V)}R$ Appendix LABEL:app:tracking_full, stability inequality derivation

### v2

@@ Abstract
Differentiable matching layers, often implemented via entropy-regularized Optimal Transport, serve as a critical approximate inference mechanism in structural prediction. However, recovering discrete permutations via annealing $\epsilon\to 0$ is notoriously unstable. We identify a fundamental mechanism for this failure: Premature Mode Collapse. By analyzing the non-normal dynamics of the Sinkhorn fixed-point map, we reveal a theoretical thermodynamic speed limit. Under standard exponential cooling, the shift in the target posterior ($O(1)$) outpaces the contraction rate of the inference operator, which degrades as $O(1/\epsilon)$. This mismatch inevitably forces the inference trajectory into spurious local basins. To address this, we propose Efficient EPH-ASC, an adaptive scheduling algorithm that monitors the stability of the inference process. By enforcing a linear stability law, we decouple expensive spectral diagnostics from the training loop, reducing overhead from $O(N^{3})$ to amortized $O(1)$. Our implementation and interactive demo are available at https://github.com/xxx0438/torch-sinkhorn-asc and https://huggingface.co/spaces/leon0923/torch-sinkhorn-asc-demo. bounded away from zero in generic training dynamics unless the feature extractor converges unrealistically fast.
@@ 3.1 Problem Setup and Sinkhorn Dynamics
See Appendix A.6 for sensitivity analysis and Appendix A.2 for spectral bounds. ∎
@@ 3.2 The Adiabatic Tracking Theorem
The proofs for Theorem 3.2 and Corollary 3.3 are provided in Appendix A.2.
@@ A.5 Directional sensitivity lower bound
A.6 Refined Sensitivity Analysis: The $O(\epsilon^{-1})$ Scaling
In this section, we provide a rigorous derivation of the sensitivity of the optimal transport plan $P^{*}_{\epsilon}$ with respect to the regularization parameter $\epsilon$. We explicitly address why the sensitivity scales as $O(\epsilon^{-1})$ rather than the $O(\epsilon^{-2})$ scaling one might expect from the Gibbs kernel derivative.
@@ A.6 Refined Sensitivity Analysis: The $O(\epsilon^{-1})$ Scaling
Let $P^{*}_{\epsilon}$ be the unique solution to the entropy-regularized optimal transport problem with cost matrix $C$. Under the non-degeneracy assumption (Assumption A.1), the Frobenius norm of the derivative of the optimal plan with respect to temperature satisfies:
$\left\|\frac{\partial P^{*}_{\epsilon}}{\partial\epsilon}\right\|_{F}=\Theta(\epsilon^{-1})\quad\text{as }\epsilon\to 0.$ (17)
@@
Recall the primal-dual relationship for the optimal plan entries on the active support:
$P^{*}_{ij}(\epsilon)=\exp\left(\frac{f_{i}(\epsilon)+g_{j}(\epsilon)-C_{ij}}{\epsilon}\right),$ (18)
where $f(\epsilon),g(\epsilon)$ are the optimal dual potentials. Taking the logarithm of both sides:
$\log P^{*}_{ij}(\epsilon)=\frac{f_{i}(\epsilon)+g_{j}(\epsilon)-C_{ij}}{\epsilon}.$ (19)
Differentiating both sides with respect to $\epsilon$ using the total derivative:
$\frac{1}{P^{*}_{ij}}\frac{\partial P^{*}_{ij}}{\partial\epsilon}=\frac{(\dot{f}_{i}+\dot{g}_{j})\epsilon-(f_{i}+g_{j}-C_{ij})}{\epsilon^{2}},$ (20)
where $\dot{f}:=\partial f/\partial\epsilon$. Rearranging terms and substituting $\frac{f_{i}+g_{j}-C_{ij}}{\epsilon}=\log P^{*}_{ij}$:
$\frac{\partial P^{*}_{ij}}{\partial\epsilon}=\frac{P^{*}_{ij}}{\epsilon}\left[(\dot{f}_{i}(\epsilon)+\dot{g}_{j}(\epsilon))-\log P^{*}_{ij}(\epsilon)\right].$ (21)
We now analyze the asymptotic magnitude of each term in the bracket as $\epsilon\to 0$:
1. The Log-Probability Term ($\log P^{*}_{ij}$): On the active support $S$ (Assumption A.1), the mass $P^{*}_{ij}$ converges to a strictly positive constant (the solution to the unregularized linear program). Thus, $\log P^{*}_{ij}=O(1)$. (For inactive entries, $P^{*}_{ij}\to 0$ exponentially fast, making the derivative vanish regardless of the polynomial factor).
2. The Dual Potential Derivatives ($\dot{f},\dot{g}$): According to the asymptotic expansion theory for entropic optimal transport [Cominetti and San Martín, 1994], the dual potentials admit a Taylor expansion of the form:
$f_{i}(\epsilon)=\phi_{i}+\epsilon\cdot h_{i}+O(\epsilon^{2}),$ (22)
where $\phi_{i}$ are the Kantorovich potentials of the unregularized problem. Differentiating this expansion with respect to $\epsilon$ yields:
$\dot{f}_{i}(\epsilon)=h_{i}+O(\epsilon)=O(1).$ (23)
The same holds for $g_{j}(\epsilon)$. Consequently, the term $(\dot{f}_{i}+\dot{g}_{j})$ is bounded by a constant $O(1)$.
Conclusion: Substituting these scalings back into Eq. (21):
$\frac{\partial P^{*}_{ij}}{\partial\epsilon}=\underbrace{\frac{P^{*}_{ij}}{\epsilon}}_{O(\epsilon^{-1})}\cdot\underbrace{\left[O(1)-O(1)\right]}_{O(1)}=O(\epsilon^{-1}).$ (24)
Thus, the sensitivity is dominated by the $1/\epsilon$ factor. ∎
Remark A.11 (Why not $O(\epsilon^{-2})$?).
It is tempting to assume the sensitivity scales as $O(\epsilon^{-2})$ by inspecting the Gibbs kernel $K_{ij}=e^{-C_{ij}/\epsilon}$, whose derivative is $\frac{C_{ij}}{\epsilon^{2}}K_{ij}$. However, this view ignores the marginal constraints. The dual potentials $f(\epsilon)$ and $g(\epsilon)$ are adaptive; they shift specifically to counteract the mass displacement caused by the changing kernel. Mathematically, this is represented by the subtraction of the $\log P^{*}_{ij}$ term in Eq. (21), which effectively cancels the higher-order singularity, reducing the divergence from $\epsilon^{-2}$ to $\epsilon^{-1}$.
@@ A.9 Raw numeric estimates and CSV
Include the CSV with columns eps, op_norm, C0_est, dS_de_norm, K1_est, K2_est, rho_mean, rho_std. When reproducing tables in the supplement please make sure the CSV file name used by matches the distributed file.
@@ B.2 Lemma: Equilibrium Error Bound
Non-normal transient amplification despite spectral contraction Theorem A.2; full proof in Appendix A.2
Effective basin shrinkage by modal condition number $\kappa(V)$ Appendix A.2, transient growth bound
Discrete-time tracking formulation for annealing Appendix B, Section “Discrete Tracking Model”
Existence of a dynamical annealing speed limit Theorem 3.2; proof and constants in Appendix B
@@
Justification of QSA drift estimation via implicit differentiation Appendix A.3; error control in Appendix B
ASC admissible drift threshold $\|\Delta_{t}\|\leq\frac{1-\rho(J_{\varepsilon})}{\kappa(V)}R$ Appendix B, stability inequality derivation